OpenStudy (anonymous):

Find and simplify the volume-to-surface-area ratio for each of the three possible ice cube shapes. You can use this chart. |Shape| S/A Formula|Volume Formula Cube | S.A=6s^2 | V=s^3 Sphere|S.A+4 pi r^2|V=4/3 pi r^3 Cylinder| S.A=2 pi r^2+2 pi rh |V= pi r^2 h please help

OpenStudy (jameshorton):

The surface area of a cube is defined as SA = 6s^2 The volume of a cube is defined as V = s^3 The surface area of 3 cubes: SA = 3*6s^3 SA = 18s^3 The volume of 3 cubes V = s^3 V = 3s^3 Volume-to-surface-area ration V:SA s^3 : 6s^2 = 3s^3 : 18s^2

OpenStudy (anonymous):

thank you so much

OpenStudy (jameshorton):

welcome